### 20 Mar 2024 v1

# M. Goncharov Maxim Goncharov

Abstract. In the paper, we introduce the notion of a Rota-Baxter operator of a non-scalar weight. As a motivation, we show that there is a natural connection between Rota-Baxter operators of this type and structures of quasitriangular Lie bialgebras on a quadratic finite-dimensional Lie algebra. Moreover, we show that some classical results on Lie bialgebra structure on simple finite-dimensional Lie algebras can be obtained from the corresponding results for Rota-Baxter operators.
Keywords: Lie algebra, Lie bialgebra, coboundary Lie bialgebra, classical Yang-Baxter equation, Rota—Baxter operator
Rota—Baxter operators for associative and commutative algebras first appeared in the paper by G. Baxter [5] as a tool for studying integral operators that appear in the theory of probability and mathematical statistics. The original notion of the Rota-Baxter operators was the following: given an associative and commutative algebra $A$, by a Baxter type operator one meant a linear map $R:A\rightarrow A$ satisfying the following quadratic equation
$R(a)R(b)=R(R(a)b+aR(b)-\theta ab),$
where $\theta\in A$ is a fixed element, $a,b\in A$ [3]. For a long period of time, this type of operators had been intensively studied in combinatorics and probability theory, mainly.
Independently, Rota—Baxter operators on Lie algebras naturally appeared in the papers of A.A. Belavin, V.G. Drinfeld [6], M.A. Semenov-Tyan-Shanskii [29] and N. Yu. Reshetikhin, M.A. Semenov-Tyan-Shanskii [28] while studying triangular and factorizable Lie bialgebras structures on simple finite-dimensional complex Lie algebras (see Section 3 for details). In this context, the definition of a Rota-Baxter operator was the following: given an arbitrary algebra $A$ over a field $F$ and $\lambda\in F$, by a Rota-Baxter operator of weight $\lambda$ one meant a linear map $R:A\rightarrow A$ satisfying
(1) $R(a)R(b)=R(R(a)b+aR(b)+\lambda ab)$
for all $a,b\in A$. Nowadays, Rota-Baxter operators of weight $\lambda$ are defined as operators satisfying (1). Operators of this type turn out to be interesting, with many interesting applications in various areas of mathematics, such as symmetric polynomials, quantum field renormalization, pre- and postalgebras, shuffle algebra, double Poisson algebras, etc. (see [1], [3], [11], [20], [25], [13]).
At the same time, the notion of a Rota-Baxter operator for groups was introduced in [21]. If $G$ is a group, then a map $B:G\mapsto G$ is called a Rota-Baxter operator on the group $G$ if for all $g,h\in G$:
$B(g)B(h)=B(gB(g)hB(g)^{-1}).$
If $(G,B)$ is a Rota-Baxter Lie group, then the tangent map of $B$ at the identity is a Rota-Baxter operator of weight 1 on the corresponding Lie algebra of the Lie group $G$. Also, it was shown that many results that are true for Rota-Baxter operators on algebras have corresponding analogs for Rota-Baxter operators on groups. And the natural question here is: what is a weight in the case of Rota-Baxter operators on groups, and how to define a Rota-Baxter operator of any weight? Considering elements of a field as a weight in this case is not natural. The same question can be asked for other algebraic structures ( rings, etc.).
In this paper, we are giving a generalization of the notion of a weight of a Rota-Baxter of an algebra. The main motivation of the paper is to provide a definition of a Rota-Baxter operator that can be used for the study of quasitriangular (not necessarily factorizable) Lie bialgebras. Instead of elements from the ground field $F$ (or elements of an associative and commutative algebra), we suggest considering centrum maps, that is, maps that commute with operators of left and right multiplications (see Section 2). This is a generalization of both notions of Rota-Baxter operators: given a scalar $\lambda\in F$ (or an element $\theta\in A$ if $A$ is an associative commutative algebra), we can consider a linear map $\varphi_{\lambda}$ (respectively, $\varphi_{\theta})$ acting as $\varphi_{\lambda}(x)=\lambda x$. From this point of view, the Rota-Baxter operator on groups from [21] can be called Rota-Baxter operators of weight $\mathrm{id}$ (where $\mathrm{id}$ is the identity map).
The paper is organized as follows: In Section 1, we give preliminary results concerning Lie bialgebras and the classical Yang-Baxter equation on Lie algebras. In Section 2, the notion of a Rota-Baxter operator of a non-scalar weight $\mu\in\mathrm{Cent}$ is introduced and studied. In Section 3, we obtain the connection between Rota-Baxter operators on a finite-dimensional Lie algebra $\mathfrak{g}$ over an arbitrary field of characteristic different from 2 and solutions of the classical Yang-Baxter equations on $\mathfrak{g}$ with $\mathfrak{g}$-invariant symmetric part (that is, with structures of triangular and quasitriangular Lie bialgebra structures on $\mathfrak{g}$). In Section 4, we use the results from Section 2 to study coboundary Lie bialgebra structures on perfect finite-dimensional quadratic Lie algebras over an arbitrary field of characteristic different from 2. Given $r\in\mathfrak{g}\wedge\mathfrak{g}$, we establish a criteria for a coalgebra $(\mathfrak{g},\delta_{r})$ to be a Lie bialgebra. Moreover, we can use the obtained results to specify whether $(\mathfrak{g},\delta_{r})$ is a triangular, quasitriangular, or factorizable Lie bialgebra. Finally, in Section 5, we show that many well-known results concerning Lie bialgebra structures on simple finite-dimensional Lie bialgebras over a field of characteristic 0 can be obtained as corollaries of the results from Sections 2-4. Finally, we show that every Lie bialgebra structure on a semisimple finite-dimensional Lie bialgebra over a field of characteristic 0 is induced by a special type of Rota-Baxter operator of some (not necessarily a scalar or even a non-degenerate) weight.
The author acknowledges the kind hospitality of Jilin University. The author is also grateful to professors Y. Sheng and R. Tang for useful discussions.
1. Preliminaries: Lie bialgebras and CYBE.
Unless otherwise specified, $F$ denotes a field of characteristic different from 2. Given vector spaces $V$ and $U$ over a field $F$, denote by $V\otimes U$ its tensor product over $F$. Define the linear mapping $\tau$ on $V\otimes V$ by $\tau(\sum\limits_{i}a_{i}\otimes b_{i})=\sum\limits_{i}b_{i}\otimes a_{i}$. If $\tau(r)=-r$, then an element $r\in V\otimes V$ is called skew-symmetric, and we will use the notation $V\wedge V$ for the subspace of skew-symmetric elements in $V\otimes V$.
Denote by $V^{*}$ the dual space of $V$. Given $f\in V^{*}$ and $v\in V$, the symbol $\langle f,v\rangle$ (or $\langle v,f\rangle$ )denotes the the linear functional $f$ evaluated on $v$ (i.e. $\langle f,v\rangle=\langle v,f\rangle=f(v)$).
Given a Lie algebra $\mathfrak{g}$, denote by $Z(\mathfrak{g})$ the center of the algebra $\mathfrak{g}$:
$Z(\mathfrak{g})=\{x\in\mathfrak{g}|\ [x,\mathfrak{g}]=0\}.$
Definition. A lie algebra $\mathfrak{g}$ is called perfect, if $[\mathfrak{g},\mathfrak{g}]=\mathfrak{g}$.
A Lie algebra $\mathfrak{g}$ acts on $\mathfrak{g}^{\otimes n}$ by
$[x_{1}\otimes x_{2}\otimes\ldots\otimes x_{n},y]=\sum\limits_{i}x_{1}\otimes\ldots\otimes[x_{i},y]\otimes\ldots\otimes x_{n}$
for all $x_{i},y\in\mathfrak{g}$.
An element $r\in L^{\otimes n}$ is called $\mathfrak{g}$-invariant (or $ad$-invariant) if $[r,y]=0$ for all $y\in L$.
Definition. Let $A$ be an arbitrary vector space over $F$. A pair $(A,\delta)$, where $\delta:A\rightarrow A\otimes A$ is a linear mapping, is called a coalgebra, while $\delta$ is a comultiplication.
Given $a\in A$, we will use the following Sweedler’s sumless notation:
$\delta(a)=a_{(1)}\otimes a_{(2)}.$
Let $(A,\delta)$ be a coalgebra. Define a multiplication on $A^{*}$ by
$fg(a)=\langle fg,a\rangle=f(a_{(1)})g(a_{(2)}),$
where $f,g\in A^{*}$, $a\in A$ and $\delta(a)=a_{(1)}\otimes a_{(2)}$. The algebra obtained is the dual algebra of the coalgebra $(A,\delta)$.
The following definition of a coalgebra related to some variety of algebras was given in [4].
Definition. Let $\mathcal{M}$ be an arbitrary variety of algebras. The pair $(A,\delta)$ is called a $\mathcal{M}$-coalgebra if $A^{*}$ belongs to $\mathcal{M}$.
The following definition was given in [10]:
Definition Let $\mathfrak{g}$ be a Lie algebra with a comultiplication $\delta$. The pair $(\mathfrak{g},\delta)$ is called a Lie bialgebra if and only if $(\mathfrak{g},\delta)$ is a Lie coalgebra and $\delta$ is a 1-cocycle, i.e., it satisfies
(2) $\delta([a,b])=[\delta(a),b]+[a,\delta(b)]$
for all $a,b\in\mathfrak{g}$.
Given a comultiplication $\delta\colon\mathfrak{g}\mapsto\mathfrak{g}\otimes\mathfrak{g}$ on a Lie algebra $\mathfrak{g}$, consider a vector space $D(\mathfrak{g})=\mathfrak{g}\oplus\mathfrak{g}^{*}$. As it was noted in [10], a pair $(\mathfrak{g},\delta)$ is a Lie bialgebra if and only if there is a (unique) structure of a Lie algebra on $D(\mathfrak{g})$ satisfying
$\mathfrak{g}$ and $\mathfrak{g}^{*}$ are subalgebras in $D(\mathfrak{g})$,
— a bilinear form $Q\colon D(\mathfrak{g})\times D(\mathfrak{g})\mapsto F$ defined as $Q(x+f,y+g)=f(y)+g(x)$ for all $x,y\in\mathfrak{g}$, $f,g\in\mathfrak{g}$ is invariant.
In this case, the product in $D(\mathfrak{g})$ is given by the formula
$[x+f,y+g]_{D}=([x,y]+\langle f,y_{(2)}\rangle y_{(1)}+\langle g,x_{(1)}\rangle x_{(2)})+([f,g]+x\bullet g+f\bullet y),$
where $\langle x\bullet g,y\rangle=\langle g\bullet y,x\rangle=g([y,x])$ for all $x,y\in\mathfrak{g}$, $f,g\in\mathfrak{g}^{*}$. The algebra $D(\mathfrak{g})$ is called the classical double [31] (or the Drinfeld double [32]) of the bialgebra $(\mathfrak{g},\delta)$.
Let $\mathfrak{g}$ be a Lie algebra and $r=\sum\limits_{i}a_{i}\otimes b_{i}\in\mathfrak{g}\otimes\mathfrak{g}$. Define a comultiplication $\Delta_{r}$ on $\mathfrak{g}$ by
$\delta_{r}(a)=[r,a]$
for all $a\in\mathfrak{g}$. It is easy to see that $\delta_{r}$ is a 1-cocycle. In [6] it was proved that $(\mathfrak{g},\Delta_{r})$ is a Lie coalgebra if and only if the following two conditions hold:
  • The symmetric part $r+\tau(r)$ of the element $r$ is $\mathfrak{g}$-invariant.
  • The element $C_{\mathfrak{g}}(r)=[r_{12},r_{13}]-[r_{23},r_{12}]+[r_{13},r_{23}]$ is $\mathfrak{g}$-invariant.
Here $[r_{12},r_{13}]=\sum\limits_{ij}[a_{i},a_{j}]\otimes b_{i}\otimes b_{j}$, $[r_{23},r_{12}]=\sum\limits_{ij}a_{i}\otimes[a_{j},b_{i}]\otimes b_{j}$, and $[r_{13},r_{23}]=\sum\limits_{ij}a_{i}\otimes a_{j}\otimes[b_{i},b_{j}]$.
The equation
(3) $C_{\mathfrak{g}}(r)=[r_{12},r_{13}]-[r_{23},r_{12}]+[r_{13},r_{23}]=0$
is called the classical Yang-Baxter equation on $\mathfrak{g}$. A solution $r\in\mathfrak{g}\otimes\mathfrak{g}$ is called a classical $r$-matrix.
Let $\mathfrak{g}$ be a finite-dimensional Lie algebra over a field $F$ and $\bar{F}$ be an arbitrary field extension of $F$ and $\delta$ be a comultiplication. Consider $\mathfrak{g}_{\overline{F}}=\mathfrak{g}\otimes_{F}\overline{F}$ as a Lie algebra over $\overline{F}$ and define
$\delta_{\overline{F}}=\delta\otimes id_{\overline{F}}:\mathfrak{g}_{\overline{F}}\mapsto(\mathfrak{g}\otimes_{F}\mathfrak{g})\otimes_{F}\overline{F}\cong(\mathfrak{g}\otimes_{F}\overline{F})\otimes_{\overline{F}}(\mathfrak{g}\otimes_{F}\overline{F})=\mathfrak{g}_{\overline{F}}\otimes_{\overline{F}}\mathfrak{g}_{\overline{F}}.$
It is straightforward to check that the classical double $D(\mathfrak{g}_{\overline{F}})$ of the bialgebra $(\mathfrak{g}_{\overline{F}},\delta_{\overline{F}})$ is isomorphic to $D(\mathfrak{g})\otimes_{F}\overline{F}$. Thus, if $(\mathfrak{g},\delta)$ is a Lie bialgebra, then $(\mathfrak{g}_{\overline{F}},\delta_{\overline{F}})$ is also a Lie bialgebra.
Definition. Let $(\mathfrak{g},\delta)$ be a Lie bialgebra. Then
(1) The pair $(\mathfrak{g},\delta)$ is called a coboundary Lie bialgebra, if $\delta=\delta_{r}$ for some $r\in\mathfrak{g}\wedge\mathfrak{g}$.
(2) The pair $(\mathfrak{g},\delta)$ is called a triangular Lie bialgebra, if $\delta=\delta_{r}$, for some $r\in\mathfrak{g}\wedge\mathfrak{g}$, where $r$ is a solution of CYBE on $\mathfrak{g}$.
(3) The pair $(\mathfrak{g},\delta)$ is called a quasitriangular Lie bialgebra, if $\delta=\delta_{r}$ for some $r\in\mathfrak{g}\otimes\mathfrak{g}$, where $r$ is a solution of CYBE on $\mathfrak{g}$.
(4) The pair $(\mathfrak{g},\delta)$ is called a factorizable Lie bialgebra, if $\delta=\delta_{r}$ for some $r\in\mathfrak{g}\otimes\mathfrak{g}$, where $r$ is a solution of CYBE on $\mathfrak{g}$ and $r+\tau(r)$ is non-degenerate (as a bilinear form on $\mathfrak{g}$) ([28]).
(5) A Lie bialgebra $(\mathfrak{g},\delta)$ over a field $F$ is called almost-factorizable, if for some field extension $\bar{F}$, the Lie bialgebra $(\mathfrak{g}\otimes_{F}\overline{F},\delta_{\overline{F}})$ is factorizable ([2]).
It is known that if $\mathfrak{g}$ is a simple complex finite-dimensional Lie algebra, then any structure of a Lie bialgebra is either triangular or factorizable, that is, is induced by a special type Rota-Baxter operator of a weight $\lambda\in\mathbb{C}$. If $\mathfrak{g}$ is an almost simple real finite-dimensional Lie algebra (that is, the complexification of $\mathfrak{g}$ is a complex simple Lie algebra), then in [2] it was noted that a Lie bialgebra structure on $\mathfrak{g}$ is either triangular, factorizable, or almost-factorizable (a similar result for reductive Lie algebras, that are extensions of finite-dimensional almost-simple real Lie algebras, follows from [12]).
2. Rota-Baxter operators of a non-scalar weights.
Let $A$ be an arbitrary algebra over a field $F$. Given $x\in A$, denote by $L_{x}$ and $R_{x}$ operators of left and right multiplications:
$L_{x}(y)=xy,\ \ R_{x}(y)=yx$
for all $y\in A$. Clearly, $L_{x},R_{x}\in\mathrm{End}(A)$. Denote by $\mathfrak{U}$ the subalgebra in $\mathrm{End}(A)$, generated by operators $L_{x}$ and $R_{x}$ for all $x\in A$: $\mathfrak{U}=alg\langle L_{x},R_{x}|x\in A\rangle$.
Definition [23]. A linear map $\mu:A\rightarrow A$ is called a centrum map in (a centralizer) of an algebra $A$ if, for any $x\in A$, $\mu\circ L_{x}=L_{x}\circ\mu$ and $\mu\circ R_{x}=R_{x}\circ\mu$. In other words, for all $x,y\in A$
$\mu(xy)=\mu(x)y=x\mu(y).$
We will denote by ${\rm Cent}(A)$ the set of all centrum maps of the algebra $A$. In other words, $\mathrm{Cent}(A)$ is the centralizer of the subalgebra $\mathfrak{U}$ in $\mathrm{End}(A)$. Clearly, ${\rm Cent}(A)$ is a subalgebra in ${\rm End}(A)$.
Note that if $A^{2}=A$, then the algebra ${\rm Cent}(A)$ is commutative. Indeed, for any $\mu,\gamma\in{\rm Cent}(A)$ and $x,y\in A$, we have
$(\varphi\circ\mu)(xy)=\varphi(x\mu(y))=\varphi(x)\mu(y)=(\mu\circ\varphi)(xy).$
Example 1. Let $A$ be an arbitrary algebra over a field $F$ and $\lambda\in F$. Define a map $\varphi_{\lambda}$ as
(4) $\varphi_{\lambda}(x)=\lambda x$
for all $x\in A$. Obviously, $\varphi_{\lambda}\in{\rm Cent(A)}$
Example 2. Let $\mathfrak{g}=sl_{2}(\mathbb{C})$ be the complex 3-dimensional special Lie algebra with the standard basis $x,h,y$ and the multiplication
$[x,y]=h,\ [h,x]=2x,\ [h,y]=-2y.$
Consider $\mathfrak{g}$ as a Lie algebra over $\mathbb{R}$. Then $\mathfrak{g}$ is a simple 6-dimensional Lie algebra and
$\mathfrak{g}=sl_{2}(\mathbb{R})\oplus isl_{2}(\mathbb{R}).$
Consider a map $\varphi_{i}$ defined as
$\varphi_{i}(a+ib)=i(a+ib)=ia-b$
for all $a,b\in sl_{2}(\mathbb{R})$. Then $\varphi_{i}\in{\rm Cent(\mathfrak{g})}$ (see [19] for details). Moreover, a map $\psi$, defined as
$\psi(a+ib)=a-ib$
for all $a,b\in sl_{2}(\mathbb{R})$, is an automorphism of $\mathfrak{g}$ (as an algebra over $\mathbb{R}$). Simple calculations show that
$\psi^{-1}\circ\varphi_{i}\circ\psi=-\varphi_{i}.$
Example 3. Suppose that an algebra $A$ splits into a direct sum of ideals $A_{i}(i=1,2,3)$: $A=A_{1}\oplus A_{2}\oplus A_{3}$. Consider a map $\mu$ defined as
$\mu(x_{1}+x_{2}+x_{3})=x_{2}+2x_{3},$
where $x_{i}\in A_{i}(i=1,2,3)$. Then $\mu$ is a centrum map with $ker(\mu)=A_{1}$.
Definition. Let $A$ be an arbitrary algebra and $\mu\in\mathrm{Cent}(A)$. A linear map $R:A\rightarrow A$ is called a Rota-Baxter operator of weight $\mu$ if for all $x,y\in A:$
(5) $R(x)R(y)=R(R(x)y+xR(y)+\mu(xy)).$
If $R$ is a Rota-Baxter operator on $A$, then we will call the pair $(A,R)$ a Rota-Baxter algebra.
Remark. If the weight $\mu$ of a Rota-Baxter operator $R$ is equal to $\varphi_{\lambda}$ (defined by (4) for some $\lambda\in F$), we will say that $R$ is a Rota-Baxter operator of weight $\lambda$ or that $R$ is a Rota-Baxter operator of a scalar weight.
Example 4. Let $A$ be the algebra of continuous functions on $\mathbb{R}$. For any $f(x)\in A$, define
[…]
Then $R$ is a Rota-Baxter operator of weight 0 (due to the integration by parts formula) [5].
Example 5. Let $d$ be an inevitable derivation of an arbitrary algebra $A$. Then $d^{-1}$ is a Rota-Baxter operator of weight 0.
The following statements generalize well-known results for Rota-Baxter operators of scalar weights to the general case.
Statement 1. Suppose an algebra $A$ splits into a direct sum of two subalgebras:$A=A_{1}\oplus A_{2}$. For any $\mu\in\mathrm{Cent}(A)$ such that $A_{1}^{2}$ and $A_{2}^{2}$ are $\mu$-invariant, a map $R$ defined as
$R(x+a)=-\mu(x)$
for all $x\in A_{1}$, $a\in A_{2}$, is a Rota-Baxter operator of weight $\mu$.
Proof. Let $x,a\in A_{1}$, $y,b\in A_{2}$. Then
$R(x+a)R(y+b)=\mu(x)\mu(y)=\mu^{2}(xy).$
On the other hand,
$R(R(x+a)(y+b)+(x+a)R(y+b)+\mu((x+a)(y+b)))\\ =R(-\mu(x)(y+b)-(x+a)\mu(y)+\mu(x)(y+b)+a\mu(y+b))\\ =R(-x\mu(y)+a\mu(b))=R(-\mu(xy)+\mu(ab))=\mu^{2}(xy).$
Statement 2. Let $R$ be a Rota-Baxter operator of weight $\mu$.
(1) If $\gamma\in{\rm Cent}(A)$, then $R^{\prime}=R\circ\gamma$ is a Rota-Baxter operator of weight $\mu\circ\gamma$.
(2) If $\varphi$ is an automorphism or an anti-automorphism of $A$, then $\varphi\circ R\circ\varphi^{-1}$ is a Rota-Baxter operator of weight $\varphi\circ\mu\circ\varphi^{-1}$.
(3) A map $R_{1}=-\mu-R$ is also a Rota-Baxter operator of the same weight $\mu$.
(4) If a linear map $Q:A\rightarrow A$ satisfies
$Q(x)Q(y)-Q(Q(x)y+xQ(y))=-\mu^{2}([x,y])$
for some $\mu\in\mathrm{Cent(A)}$ and all $x,y\in A$, then a map $R=\frac{1}{2}(Q-\mu)$ is a Rota-Baxter operator of weight $\mu$.
Proof. (1) For any $x,y\in A$, we have
$R^{\prime}(x)R^{\prime}(y)=R(\gamma(x))R(\gamma(y))=R(R(\gamma(x))\gamma(y)+\gamma(x)R(\gamma(y))+\mu(\gamma(x)\gamma(y)))\\ =R(\gamma(R^{\prime}(x)y+xR^{\prime}(y)+\gamma(x)\mu(\gamma(y)))=R^{\prime}(R^{\prime}(x)y+yR^{\prime}(x)+(\mu\circ\gamma)(xy)).$
(2) Similar to (1).
(3) For any $x,y\in\mathfrak{g}$, we have
$R_{1}(x)R_{1}(y)=(\mu+R)(x)(\mu+R)(y)\\ =\mu^{2}(xy)+\mu(R(x)y+xR(y))+R(x)R(y).$
On the other hand,
$R_{1}(R_{1}(x)y+xR_{1}(y)+\mu(xy))\\ =R_{1}(-\mu(x)y-R(x)y-x\mu(y)-xR(y)+\mu(xy))\\ =(\mu+R)(R(x)y+xR(y)+\mu(xy))=\mu(R(x)y+xR(y)+\mu(xy))+R(x)R(y).$
(4) similar to (3). The statement is proved.
Remark. Results from Statement 2 were obtained in [5] in the case when $\mu=\varphi_{\lambda}$, $\lambda\in F$.
If $R:\mathfrak{g}\rightarrow\mathfrak{g}$ is a Rota-Baxter operator of a scalar weight $\lambda$, a classical result says that a pair $(\mathfrak{g},[,]_{R})$, where
$[x,y]_{R}=[R(x),y]+[x,R(y)]+\lambda[x,y],\ x,y\in\mathfrak{g},$
is also a Lie algebra [29]. a Similar technique allows one to obtain a similar result in the case when the weight of $R$ is an arbitrary $\mu\in\mathrm{Cent}(\mathfrak{g})$ (and arbitrary varieties of algebras). Nevertheless, we will use the following construction that we find interesting and that we will use in the future. This construction (with a little difference) originally appears in [13] for Rota-Baxter operators of scalar weights.
Let $A$ be an algebra over an arbitrary field $F$, $R\colon A\mapsto A$ be a linear map, and $\mu\in\mathrm{Cent}(A)$. Consider a direct sum of vector spaces $D_{R,\mu}(A)=A\oplus\bar{A}$, where $\bar{A}$ is an isomorphic copy of $A$. Define a product on $D_{R,\mu}(A)$ as follows:
$(a+\bar{b})*(x+\bar{y})=ax+R(ay)-aR(y)+R(bx)-R(b)x\\ +\overline{ay+bx-R(b)y-bR(y)-\mu(by)},$
$a,b,x,y\in A$.
Obviously, $A$ is a subalgebra in $D_{R,\mu}(A)$.
For any $x\in A$, define
(6) $i(x)=\bar{x}+(\mu(x)+R(x))\in D_{R,\mu}(A).$
Define
(7) $I(A)=\{i(x)\mid x\in A\}$
Clearly, $I(A)$ is a subspace of $D_{R,\mu}(A)$ isomorphic to $A$ as a vector space, and $I(A)\cap A=0$. Moreover,
$D_{R,\mu}(A)=A\oplus I(A)$
as a vector space. Also, for any $x,y\in A$, we have
(8) $i(x)*y=(\bar{x}+\mu(x)+R(x))*y\\ =\overline{xy}+R(xy)-R(x)y+\mu(x)y+R(x)y=i(xy),$
that is, $A*I(A)\subset I(A)$. Similarly, $y*i(x)=i(yx)$ and $I(A)*A\in I(A)$.
Statement 3. $I(A)$ is an ideal in $D_{R,\mu}(A)$ if and only if $R$ is a Rota-Baxter operator of weight $\mu$.
Proof. In light of (8), we need to check when $i(x)*\overline{y}\in I(A)$ and $\overline{y}*i(x)\in I(A)$. Consider $x,y\in A$, then
$i(x)*\bar{y}=(\bar{x}+\mu(x)+R(x))*\bar{y}\\ =\overline{-\mu(xy)-xR(y)-R(x)y+\mu(x)y+R(x)y}\\ +R(\mu(x)y)-\mu(x)R(y)+R(R(x)y)-R(x)R(y)\\ =\overline{-xR(y)}-\mu(xR(y))-R(xR(y))+R(\mu(xy)+R(x)y+xR(y))-R(x)R(y)\\ =i(-xR(y))+R(\mu(xy)+R(x)y+xR(y))-R(x)R(y).$
Since $I(A)\cap A=0$, we obtain that $i(x)*\overline{y}$ belongs to $I(A)$ for any $x,y\in A$ if and only if $R$ is a Rota-Baxter operator of weight $\mu$. Similarly, $\bar{y}*i(x)\in I(A)$ if and only if $R$ is a Rota-Baxter operator of weight $\mu$. The statement is proved.
Note that if $R$ is a Rota-Baxter operator of weight $\mu$, we have for all $x,y\in A$
(9) $i(x)*i(y)=i(x)*(\bar{y}+\mu(y)+R(y))\\ =-i(xR(y))+i(\mu(xy+xR(y))=i(\mu(xy)).$
If $\mu$ is invertible, then a map $i_{\mu}\colon x\mapsto i(\mu^{-1}(x))$ is an isomorphism of algebras $A$ and $I(A)$. If $\mu=0$, then $I(A)^{2}=0$.
Corollary 1. Let $A$ be an algebra over a field $F$ and $R:A\rightarrow A$ be a Rota-Baxter operator of weight $\mu\in\mathrm{Cent}(A)$. Then a non-associative polynomial $f(t_{1},\ldots,t_{n})$ in non-commuting variables $t_{1},\ldots,t_{n}$ is an identity in $A$ if and only if $f(t_{1},\ldots,t_{n})$ is an identity in $(D_{R,\mu}(A),*)$. In particular, algebras $(D_{\mu,R(A)},*)$ and $A$ belong to the same variety of algebras.
Proof. Recall that $D_{R,\mu}(A)=A\oplus I(A)$. Let $f(t_{1}\ldots t_{n})$ be a non-associative polynomial in non-commuting variables $t_{1},\ldots t_{n}$. Suppose that $f(t_{1},\ldots,t_{n})$ is an identity in $A$, that is, for all $a_{1},\ldots,a_{n}\in A$: $f(a_{1},\ldots,a_{n})=0$. Consider homogeneous elements $h_{i}\in A\cup I(A)$, $i=1,\ldots,n$. If for all $i=1,\ldots n$ $h_{i}\in A$, then $f(h_{1},\ldots,h_{n})=0$ by the assumption. In general, we have that $f(h_{1},\ldots,h_{n})=i(\mu^{k}(f(x_{1},\ldots,x_{n}))=0$, where $k$ is a number of elements $h_{1},\ldots h_{n}$ from $I(A)$ and $x_{s}=h_{s}$ if $h_{s}\in A$ and $x_{s}=i^{-1}(h_{s})$ if $h_{s}\in I(A)$.
Since $A$ is a subalgebra in $D_{R,\mu}(A)$, $A$ satisfy any identity of the algebra $D_{R,\mu}(A)$. The corollary is proved.
Corollary 2. Let $\mathcal{M}$ be a variety of algebras, and $A\in\mathcal{M}$. Suppose that $R$ is a Rota-Baxter operator of weight $\mu\in\mathrm{Cent}(A)$. Define a new product $\cdot_{R}$ by
$x\cdot_{R}y=R(x)y+xR(y)+\mu(xy),$
where $x,y\in A$. Then the algebra $(A,\cdot_{R})$ lies in the variety generated by the algebra $A$. In particular, $(A,\cdot_{R})$ is an algebra from $M$.
Corollary 3. Let $A$ be an algebra over an arbitrary field $F$ and $R:A\rightarrow A$ be a Rota-Baxter operator of weight $\mu\in\mathrm{Cent}(A)$. Then
(1) If $I(A)$ is the ideal in $D_{R,\mu}(A)$ defined in (7), then $I(A)^{2}=0$ if and only if $R$ is a Rota-Baxter operator of weight 0.
(2) If $\mu$ is an invertible map, then $D_{R,\mu}(A)$ is isomorphic to the direct sum of ideals $A\oplus A$.
Proof. From (9), it follows that $I(A)$ is an abelian ideal if and only if $\mu(xy)=0$ for all $x,y\in A$. Thus, the summand $\mu(xy)$ may be omitted and $R$ is a Rota-Baxter operator of weight 0.
Consider the second statement. If $\mu$ is invertible, then a map $\phi:A\mapsto I(A)$ defined as
$\psi(x)=i(\mu^{-1}(x))$
is an isomorphism of algebras. Consider $J(A)=\{j(x)=\mu(x)-i(x)|x\in A\}$. From the definition of $I(A)$, we get that $I(A)\cap A=0$. Therefore, $I(A)\cap J(A)=0$. For any $x\in A$
$x=i(\mu^{-1}(x))+j(\mu^{-1}(x)).$
Thus, $D_{R}(A)=I(A)\oplus J(A)$. From (8) we get that
$j(x)*y=(\mu(x)-i(x))*y=\mu(xy)-i(xy)=j(xy)\in J(A)$
for any $x,y\in A$. From (8) and (9), we obtain
$j(x)*i(y)=(\mu(x)-i(x))*i(y)=i(\mu(xy))-i(\mu(xy))=0$
for any $x,y\in A$. Simirarely, $y*j(x)\in J(A)$ and $i(y)j(x)=0$. Therefore, $J(A)$ is an ideal in $D_{R,\mu}(A)$, and $I(A)*J(A)=J(A)*I(A)=0$. Finally, for any $x,y\in A$,
$j(x)*j(y)=(\mu(x)-i(x))*j(y)=j(\mu(xy)).$
Thus, algebras $I(A)$, $J(A)$ and $A$ are isomorphic. The corollary is proved. We will also need the following
Proposition 1. Let $A$ be a simple algebra, $R\in\mathrm{End}(A)$ and $\mu\in\mathrm{End}(A)$. If $I$ is a proper ideal in $D_{R,\mu}$, then $I$ is isomorphic to $A$ as a vector space, and $D_{R,\mu}(A)=A\oplus I$ (a semidirect sum).
Proof. Let $I$ be a proper ideal in $D_{R,\mu}(A)$. Then the intersection $I_{1}=I\cap A$ is an ideal in $A$. Therefore, $I_{1}=A$ or $I_{1}=0$.
If $I_{1}=A$, then for any $a,b\in A$, we have that
$a*\overline{b}=\overline{ab}+x$
for some $x\in A$. Since $A^{2}=A$ and $A\subset I$, we get that $\overline{A}\subset I$. Thus, $I=D_{R,\mu}(A)$, a contradiction.
Therefore, $A\cap I=0$. Let $W=\{a\in A|\ \overline{a}+b\in I\ \text{for some}\ b\in A\}$. Consider $a\in W$ and $x\in A$. Suppose that $\overline{a}+b\in I$, where $b\in A$. Then
$(\overline{a}+b)*x=\overline{ax}+y$
for some $y\in A$. Thus, $W$ is an ideal in $A$. If $W=0$, then $I=0$. Therefore, $W=A$. If for some $a\in A$ there are $b_{1},b_{2}\in A$ such that $\overline{a}+b_{1}\in I$ and $\overline{a}+b_{2}\in I$, then $b_{1}-b_{2}\in I\cap A=0$. Thus, there is a linear function $\gamma\in\mathrm{End}(A)$ such that $I=\{\overline{a}+\gamma(a)|a\in A\}$. A map $\varphi:A\rightarrow I$ acting as
$\varphi(a)=\overline{a}+\gamma(a)$
for $a\in A$, is an isomorphism of vector spaces.
Since $\overline{A}\subset A\oplus I$, we have that $D_{R,\mu}(A)=A\oplus I$. The proposition is proved.
3. Rota-Baxter operators and CYBE
Rota-Baxter operators (of a scalar weight) naturally appear in the paper of M.A. Semenov-Tyan-Shanskii [29]. In this paper (titled "What a classical $r$-matrix really is", a "right" version of the notion of a classical $r$ matrix was suggested. Instead of considering the notion of Lie bialgebra, he introduced the notion of a double Lie algebra (which should not be confused with the notion of the double Lie algebra from [30] and [13]).
Definition ([29]). Given a Lie algebra $\mathfrak{g}$, a linear map $R\in\textrm{End}(\mathfrak{g})$ is called a classical $R$-matrix if the space $\mathfrak{g}$ with a new product $[\cdot,\cdot]_{R}$, defined as
$[x,y]_{R}=[R(x),y]+[x,R(y)],$
where $x,y\in\mathfrak{g}$, is a Lie algebra.
Remark. In [29], it was noticed that the notion of the double Lie algebra coincides with the notion of the Lie bialgebra if and only if the map $R$ is skew-symmetric with respect to some non-degenerate bilinear quadratic form, see Section 3 for details.
Given a map $R:\mathfrak{g}\mapsto\mathfrak{g}$, define a bilinear map $\theta_{R}$ as
$\theta_{R}(x,y)=[R(x),R(y)]-R([R(x),y]+[x,R(y)]).$
Proposition 2 ([29]) $R$ is a classical R-matrix if and only if
(10) $[\theta_{R}(x,y),z]+[\theta_{R}(y,z),x]+[\theta_{R}(z,x),y]=0,\ \forall x,y,z\in\mathfrak{g}.$
M.A Semenov-Tan-Shanskii suggested the following two conditions, each of which implies (10):
i) $\theta_{R}(x,y)=0$.
In this case, $R$ is a Rota-Baxter operator of weight zero. In [29], this equation is also called the classical Yang–Baxter equation.
ii) $\theta_{R}(x,y)=-\lambda^{2}[x,y]$, $\lambda\neq 0$.
This equation is called the modified classical Yang–Baxter equation (mCYBE). It can be written as
(11) $[R(x),R(y)]=R([R(x),y]+[x,R(y)])-\lambda^{2}[x,y].$
The parameter was chosen so the map $R=\lambda\textrm{id}$ is a solution of (11).
Conditions (i) and (ii) are easier to work with since they are bilinear rather than trilinear.
It turns out that a map $R\in\textrm{End}(\mathfrak{g})$ is a solution of mCYBE (11) if and only if a map
$B=\frac{1}{2}(R-\lambda\textrm{id})$
is a Rota–Baxter operator of weight $\lambda$ on $\mathfrak{g}$.
Definition Let $\mathfrak{g}$ be a Lie algebra. A bilinear form $\omega$ on $\mathfrak{g}$ is called invariant if for all $a,b,c\in\mathfrak{g}$:
$\omega([a,b],c)=\omega(a,[b,c]).$
Definition. Let $\mathfrak{g}$ be a Lie algebra, and $\omega$ is a bilinear invariant non-degenerate form on $\mathfrak{g}$. Then the pair $(\mathfrak{g},\omega)$ is called a quadratic Lie algebra.
Notation. Let $(\mathfrak{g},\omega)$ be a quadratic Lie algebra and $R:\mathfrak{g}\rightarrow\mathfrak{g}$ be a linear map. Then by $R^{*}$ we will denote the adjoint to $R$ with respect to the form $\omega$ map defined as
$\omega(R(x),y)=\omega(x,R^{*}(y))$
for all $x,y\in\mathfrak{g}$.
Remark. Let $\mathfrak{g}$ be a perfect Lie algebra, $\omega$ be a non-degenerate associative bilinear form on $\mathfrak{g}$, and $\mu\in\mathrm{Cent}(\mathfrak{g})$. Then $\mu^{*}=\mu$. Indeed, for all $x,y,z\in\mathfrak{g}$, we have
$\omega(\mu([x,y]),z)=\omega([z,\mu(y)],z)=\omega(z,[\mu(y),z])=\omega(z,[y,\mu(z)])=\omega([z,y],\mu(z)).$
Since $[\mathfrak{g},\mathfrak{g}]=\mathfrak{g}$, we conclude that $\mu^{*}=\mu$.
If $(\mathfrak{g},\omega)$ is a quadratic finite-dimensional Lie algebra, then there is a natural isomorphism :*$*{}^{*}\colon\mathfrak{g}\mapsto\mathfrak{g}^{*}$ between the vector space $\mathfrak{g}$ and the dual space $\mathfrak{g}^{*}$ defined as $x^{*}(y)=\omega(x,y)$. This, in turn, gives a natural isomorphism between $\textrm{End}(\mathfrak{g})\cong\mathfrak{g}\otimes\mathfrak{g}^{*}$ and $\mathfrak{g}\otimes\mathfrak{g}$. This allows us to consider an element $r=\sum a_{i}\otimes b_{i}\in\mathfrak{g}\otimes\mathfrak{g}$ as a linear map $R_{r}$ acting as
(12) $R_{r}(x)=\sum\omega(a_{i},x)b_{i}.$
Remark. Note that the adjoint map $R_{r}$ with respect to the form $\omega$ is defined by
$R^{*}_{r}(x)=\sum\omega(b_{i},x)a_{i}$
for all $x\in\mathfrak{g}$.
Statement 4 ([29]). Let $(\mathfrak{g},\omega)$ be a finite-dimensional quadratic Lie algebra. An element $r\in\mathfrak{g}\otimes\mathfrak{g}$ is a skew-symmetric solution of the classical Yang-Baxter equation (3) if and only if the corresponding map $R_{r}$ is a skew-symmetric Rota-Baxter operator of weight 0.
In other words, skew-symmetric solutions of the classical Yang-Baxter equation (3) are in one-to-one correspondence with skew-symmetric Rota-Baxter operators in the case of quadratic Lie algebras.
Consider a structure of a factorizable Lie bialgebra $(\mathfrak{g},\delta_{r})$ on a Lie algebra $\mathfrak{g}$. Since $I=r+\tau(r)$ is non-degenerate, it defines an isomorphism of vector spaces $I:\mathfrak{g}^{*}\mapsto\mathfrak{g}$: $I(f)=(f\otimes id)(r)$. In [28] it was noted that a form $\beta$ defined as
(13) $\beta(x,y)=\langle I^{-1}(x),y\rangle$
for all $x,y\in\mathfrak{g}$ is a nondegenerate invariant form on $\mathfrak{g}$. Moreover, in [7] it was proved that the corresponding to $r$ and $\beta$ linear map $R_{r}$ is a Rota-Baxter operator of weight 1 satisfying
(14) $R_{r}+R_{r}^{*}+id=0.$
In other words, there is a one-to-one correspondence between structures of a factorizable Lie bialgebras on a finite-dimensional Lie algebra $\mathfrak{g}$ and triples $(\mathfrak{g},\beta,R)$, where $\beta$ is a non-degenerate invariant form on $\mathfrak{g}$ and $R$ is a Rota-Baxter operator of weight 1 satisfying (14).
Given a finite-dimensional Lie algebra $\mathfrak{g}$, there may be more than one non-degenerate invariant form on $\mathfrak{g}$ (even if $\mathfrak{g}$ is simple). We can start with a quadratic finite-dimensional coboundary Lie bialgebra $(\mathfrak{g},\delta_{r})$, $r\in\mathfrak{g}\otimes\mathfrak{g}$ and ask when $(\mathfrak{g},\delta_{r})$ is factorizable. We obtain the following version of the result from [28].
Theorem 1. Let $\mathfrak{g}$ be a finite-dimensional Lie algebra over a field $F$ and $\omega$ be a non-degenerate bilinear invariant form on $\mathfrak{g}$. Let $r=\sum\limits_{i}a_{i}\otimes b_{i}\in\mathfrak{g}\otimes\mathfrak{g}$ and $R_{r}$ be the corresponding map defined as in (12). Then
(1) $r$ is a solution of CYBE with $\mathfrak{g}$-invariant symmetric part $r+\tau(r)$ if and only if the map $R_{r}$ is a Rota-Baxter operator of a weight $\mu\in\mathrm{Cent}(\mathfrak{g})$ satisfying
(15) $R_{r}+R_{r}^{*}+\mu=0,$
where $R^{*}$ is the adjoint to $R$ with respect to the form $\omega$ map.
(2) $r$ is a solution of CYBE with $\mathfrak{g}$-invariant and non-degenerate symmetric part $r+\tau(r)$ if and only if the map $R_{r}$ is a Rota-Baxter operator of a non-degenerate weight $\mu\in\mathrm{Cent}(\mathfrak{g})$ satisfying (15).
Proof. For $x,y\in\mathfrak{g}$, consider a map $\psi_{x,y}:\mathfrak{g}\otimes\mathfrak{g}\otimes\mathfrak{g}\rightarrow\mathfrak{g}$ defined as
$\psi_{x,y}(a\otimes b\otimes c)=\omega(x,a)\omega(y,b)c$
for all $a,b,c\in\mathfrak{g}$.
Since $\omega$ is non-degenerate, $C_{\mathfrak{g}}(r)=0$ if and only if $\psi_{x,y}(C_{\mathfrak{g}}(r))=0$ for all $x,y\in\mathfrak{g}$. Direct computations show
$\psi_{x,y}(C_{\mathfrak{g}}(r))=\sum\limits_{i,j}\omega([a_{i},a_{j}],x)\omega(b_{i},y)b_{j}-\omega(a_{i},x)\omega([a_{j},b_{i}],y)b_{j}+\omega(a_{i},x)\omega(a_{j},y)[b_{i},b_{j}]=$
$=R_{r}([x,R^{*}_{r}(y)]-[R(x),y])+[R_{r}(x),R_{r}(y)].$
That is, $r$ is a solution of CYBE if and only if for all $x,y\in\mathfrak{g}$
(16) $R_{r}([x,R^{*}_{r}(y)]-[R(x),y])+[R_{r}(x),R_{r}(y)]=0.$
Similarly, $r+\tau(r)$ is $\mathfrak{g}$-invariant if and only if for all $x,y\in\mathfrak{g}$
$[(R_{r}+R_{r}^{*})(x),y]=(R_{r}+R_{r}^{*})([x,y]),$
that is, $R_{r}+R_{r}^{*}$ satisfies (15) for some $\mu\in\mathrm{Cent}(\mathfrak{g})$. Note that (16) and (15) are equivalent to the fact that $R_{r}$ is a Rota-Baxter operator of weight $\mu$ satisfying (15). Finally, $r+\tau(r)$ is non-degenerate if and only if $\mu=-R_{r}-R^{*}_{r}$ is a non-degenerate map. The theorem is proved.
Statement 5. Let $\mathfrak{g}$ — be a Lie algebra over a field $F$ with $Z(\mathfrak{g})=0$, $\omega:\mathfrak{g}\times\mathfrak{g}\mapsto F$ be a bilinear invariant non—degenerate form on $\mathfrak{g}$. Suppose that $\theta:\mathfrak{g}\times\mathfrak{g}\mapsto\mathfrak{g}$ is a bilinear map satisfying the following conditions
(17) $\theta(x,y)=-\theta(y,x),$
(18) $\omega(\theta(x,y),z)=\omega(x,\theta(y,z))$
(19) $[\theta(x,y),z]=\theta([x,z],y)+\theta(x,[y,z]),$
for all $x,y,z\in\mathfrak{g}$. Then there is a linear map $\nu:[\mathfrak{g},\mathfrak{g}]\rightarrow\mathfrak{g}$ such that $\theta(x,y)=\nu([x,y])$ for all $x,y\in\mathfrak{g}$. Moreover, for all $x,y,z\in\mathfrak{g}$ the map $\nu$ satisfies
(20) $\nu([[x,y],z])=[\nu([x,y]),z].$
Proof.
Let $x,y,z\in\mathfrak{g}$. From (17) and (19) it follows that
(21) $[\theta(x,y),z]=\theta([x,z],y)+\theta(x,[y,z])=[x,\theta(z,y)]-\theta(z,[x,y])+[y,\theta(x,z)]-\theta([y,x],z)\\ =[x,\theta(z,y)]+[\theta(z,x),y]-2\theta(z,[x,y]).$
Therefore, for all $x,y,z,t\in\mathfrak{g}$:
(22) $\omega\left([\theta(x,y),z],t\right)=\omega\left([x,\theta(z,y)]+[\theta(z,x),y]-2\theta(z,[x,y]),t\right)\\ =\omega(z,\theta(y,[t,x])+\theta(x,[y,t]))-2\omega(\theta(z,[x,y]),t)=\omega(z,[t,\theta(y,x)])-2\omega(\theta(z,[x,y]),t)\\ =\omega([z,\theta(x,y)]-2\theta(z,[x,y]),t).$
Since the form $\nu$ is non-degenerate, we have
(23) $[\theta(x,y),z]=\theta([x,y],z).$
Consider $x=\sum\limits_{i}[x_{i},y_{i}]\in[\mathfrak{g},\mathfrak{g}]$, where $x_{i},y_{i}\in\mathfrak{g}$. Define
$\nu(x)=\sum\limits_{i}\theta(x_{i},y_{i}).$
First, we need to prove that $\nu$ is well-defined. Indeed, if $\sum[x_{i},y_{i}]=\sum[z_{i},q_{i}]$ for some $x,y,z,q\in\mathfrak{g}$, then we can use (23) and note that for any $z\in\mathfrak{g}$ we have:
$\sum[\theta(x_{i},y_{i}),z]=\theta\left(\sum[x_{i},y_{i}],z\right)=\theta\left(\sum[z_{i},q_{i}],z\right)=\sum[\theta(z_{i},q_{i}),z].$
Since the center of the algebra $\mathfrak{g}$ is equal to 0, we get that $\sum\theta(x_{i},y_{i})=\sum\theta(z_{i},q_{i})$. Thus, $\nu$ is well-defined. Obviously, $\nu$ is linear.
Consider $x,y,z\in\mathfrak{g}$. Then
$\nu([[x,y],z])=\theta([x,y],z)=[\theta(x,y),z]=[\nu([x,y]),z].$
The statement is proved.
Corollary 4. Let $\mathfrak{g}$ be a semisimple finite-dimensional Lie algebra over a field of characteristic 0, and let $\chi:\mathfrak{g}\times\mathfrak{g}\mapsto F$ be the Killing form of the algebra $\mathfrak{g}$. Suppose that $\theta:\mathfrak{g}\times\mathfrak{g}\mapsto\mathfrak{g}$ is a bilinear map satisfying conditions (17)-(19). Then there is a map $\nu:\mathfrak{g}\rightarrow\mathfrak{g}$ from the centroid of the algebra $\mathfrak{g}$ such that for all $x,y\in\mathfrak{g}$:
$\theta(x,y)=\nu([x,y])=[\nu(x),y]=[x,\nu(y)].$
Proof. In this case, $[\mathfrak{g},\mathfrak{g}]=\mathfrak{g}$ and $Z(\mathfrak{g})=0$. The rest follows from Statement 5.
4. Coboundary Lie bialgebra structures and Rota-Baxter operators.
Let $(\mathfrak{g},\omega)$ be a quadratic Lie algebra over a field different from 2, $r=\sum\limits_{i}a_{i}\otimes b_{i}\in\mathfrak{g}\wedge\mathfrak{g}$. Define a comultiplication $\delta_{r}$ as
$\delta_{r}(x)=[r,x]=\sum\limits_{i}[a_{i},x]\otimes b_{i}+a_{i}\otimes[b_{i},x]$
for all $x\in\mathfrak{g}$.
Consider a map $R_{r}:\mathfrak{g}\rightarrow\mathfrak{g}$ defined as
$R_{r}(x)=\sum\limits_{i}\omega(a_{i},x)b_{i}$
for all $x\in\mathfrak{g}$. Since $r$ is skew-symmetric, $R^{*}=-R$. Recall that given $x\in\mathfrak{g}$, by $x^{*}$ we will denote the element from the dual space $\mathfrak{g}^{*}$ defined as $\langle x^{*},y\rangle=\omega(x,y)$ for all $y\in\mathfrak{g}$.
There are several questions that one can ask: when a pair $(\mathfrak{g},\delta_{r})$ is a Lie bialgebra? If $(\mathfrak{g},\delta_{r})$ is a Lie bialgebra, is it a triangular, factorizable, or almost factorizable Lie bialgebra? What can we say about the classical double of $(\mathfrak{g},\delta_{r})$? In this section, we will consider these questions from the point of view of the map $R_{r}$.
Proposition 3. Let $(\mathfrak{g},\omega)$ be an arbitrary finite-dimentional quadratic Lie algebra, and let $R\colon\mathfrak{g}\rightarrow\mathfrak{g}$ be a linear operator satisfying $R+R^{*}+\mu=0$ for some $\mu\in\mathrm{Cent}(\mathfrak{g})$ and $r\in\mathfrak{g}\otimes\mathfrak{g}$ be the corresponding tensor defined by (12). Then the multiplication in the dual algebra of the coalgebra $(\mathfrak{g},\delta_{r})$ is given by
$[a^{*},b^{*}]=(-[R(a),b]-[a,R(b)]-\mu([a,b]))^{*}.$
for all $a,b\in\mathfrak{g}$. Moreover, the classical double $D(g)=\mathfrak{g}\oplus\mathfrak{g}^{*}$ of the pair $(\mathfrak{g},\delta_{r})$ (with the product defined by $[\cdot,\cdot]_{D}$) is isomorphic to the algebra $D_{R,\mu}(\mathfrak{g})=\mathfrak{g}\oplus\overline{\mathfrak{g}}$ with a product $[\cdot,\cdot]_{R}$ given by
$[x+\bar{y},z+\bar{t}]_{R}=[x,z]+R([x,t])-[x,R(t)]+R([y,z])-[R(y),z]\\ +\overline{[x,t]+[y,z]-[R(y),t]-[y,R(t)]-\mu([y,t])}$
for all $x,y,z,t\in\mathfrak{g}$.
Proof. Let $r=\sum a_{i}\otimes b_{i}$. Then $\delta_{r}(t)=\sum[a_{i},t]\otimes b_{i}+a_{i}\otimes[b_{i},t]$ for all $t\in\mathfrak{g}$. Consider a map $\varphi:D(\mathfrak{g})\rightarrow D_{r,\mu}(\mathfrak{g})$ defined as
$\varphi(x+y^{*})=x+\overline{y}$
for all $x,y\in\mathfrak{g}$. Obviously, $\varphi$ is an isomorphism of vector spaces, and $\varphi([x,y]_{D})=[\varphi(x),\varphi(y)]_{R}$ for all $x,y\in\mathfrak{g}$.
For any $t\in\mathfrak{g}$ we have
$\langle[x^{*},y^{*}]_{D},t\rangle=\omega(x,t_{(1)})\omega(y,t_{(2)})=\sum\omega(x,[a_{i},t])\omega(y,b_{i})+\omega(x,a_{i})\omega(y,[b_{i},t])\\ =\sum\omega([t,x],a_{i})\omega(y,b_{i})+\omega(x,a_{i})\omega([t,y],b_{i})=\omega([t,x],R^{*}(y))+\omega([t,y],R(x))\\ =\omega(t,-[x,R(y)]-[x,\mu(y)]-[R(x),y])=\langle(-[R(x),y]-[x,R(y)]-\mu([x,y]))^{*},t\rangle.$
Hence, $[x^{*},y^{*}]_{D}=(-[R(x),y]-[x,R(y)]-\mu([x,y]))^{*}$ and $\varphi([x^{*},y^{*}]_{D})=[\varphi(x^{*}),\varphi(y^{*})]_{R}$.
Consider $\varphi([x,y^{*}]_{D})$ for some $x,y\in\mathfrak{g}$. For any $t\in\mathfrak{g}$, using the associativity of the form $Q$, we have the following equations:
(24) $Q([x,y^{*}]_{D},t^{*})=Q(x,[y^{*},t^{*}]_{D})=-\omega(x,[R(y),t]+[y,R(t)]+\mu([y,t]))\\ =\omega(-[x,R(y)],t)+\omega(x,[y,R^{*}(t)])=\omega(-[x,R(t)]+R([x,y]),t)$
(25) $Q([x,y^{*}]_{D},t)=Q(y^{*},[t,x])=\omega(y,[t,x])=\omega([x,y],t)=Q([x,y]^{*},t).$
From (24) and (25) it follows that
$[x,y^{*}]_{D}=R([x,y])-[x,R(y)]+[x,y]^{*}.$
Therefore, $\varphi([x,y^{*}]_{D})=[\varphi(x),\varphi(y^{*})]_{R}$, and $\varphi$ is an isomorphism of algebras. The proposition is proved.
From Theorem 1, Proposition 3, and Corollary 3, it follows
Corollary 5.
(1) If $(\mathfrak{g},\delta_{r})$ is a triangular Lie bialgebra, then the classical double is a semidirect sum $D(\mathfrak{g})=\mathfrak{g}\oplus D$, where $D$ is an abelian ideal in $D(\mathfrak{g})$.
(2) If $(\mathfrak{g},\delta_{r})$ is a factorizable Lie bialgebra, then the classical double $D(\mathfrak{g})$ is isomorphic to the direct sum of two ideals $\mathfrak{g}\oplus\mathfrak{g}$.
Definition. Given a Lie algebra $\mathfrak{g}$ and a linear map $R:\mathfrak{g}\rightarrow\mathfrak{g}$, define a linear map $\theta:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$ as
(26) $\theta(x,y)=[R(x),R(y)]-R_{r}([R(x),y]+[x,R(y)]$
for all $x,y\in\mathfrak{g}$.
Proposition 4. Let $(\mathfrak{g},\omega)$ be a quadratic Lie algebra, $r\in\mathfrak{g}\wedge\mathfrak{g}$, $R_{r}$ be the corresponding map, and $\theta:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$ be a bilinear map defined by (26). Then for all $x,y,z\in\mathfrak{g}$
(27) $\theta(x,y)=-\theta(y,x),$
(28) $\omega(\theta(x,y),z)=\omega(x,\theta(y,z)).$
Proof. The equality (27) is obvious. Let $x,y,z\in\mathfrak{g}$. Then
$\omega(\theta(x,y),z)=\omega([R_{r}(x),R_{r}(y)]-R_{r}([R_{r}(x),y]+[x,R_{r}(y)]),z)\\ =\omega(R(x),[R(y),z]-[y,R^{*}_{r}(z)])-\omega(x,[R_{r}(y),R^{*}(z)])\\ =\omega(x,[R_{r}(y),R_{r}(z)]-R_{r}([R_{r}(y),z]+[y,R_{r}(z)])=\omega(x,\theta(y,z)).$
Proposition 5. Let $(\mathfrak{g},\omega)$ be a quadratic Lie algebra, $r\in\mathfrak{g}\wedge\mathfrak{g}$, $R_{r}$ be the corresponding map, and $\theta:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$ be a bilinear map defined by (26). Then the pair $(\mathfrak{g},\delta_{r})$ is a Lie bialgebra if and only if the following condition holds
(29) $[\theta(x,y),z]=\theta([x,z],y)+\theta(x,[y,z])$
for all $x,y,z\in\mathfrak{g}.$
Proof. Since $\delta_{r}$ satisfies (2) for any $r\in\mathfrak{g}\otimes\mathfrak{g}$, we need to check when the pair $(\mathfrak{g},\delta_{r})$ is a Lie coalgebra. From Proposition 3, we know that $(\mathfrak{g},\delta_{r})$ is a Lie coalgebra if and only if the pair $(\mathfrak{g},[\cdot,\cdot])$, where the product is defined as
$[x,y]_{R_{r}}=[R_{r}(x),y]+[x,R_{r}(y)]$
for all $x,y\in\mathfrak{g}$, is a Lie algebra. From Proposition 2, it follows that $(\mathfrak{g},[\cdot,\cdot]_{R_{r}})$ is a Lie algebra if and only if the bilinear map $\theta$ defined in (26) satisfies
(30) $[\theta(x,y),z]+[\theta(y,z),x]+[\theta(z,x),y]=0$
for all $x,y,z\in\mathfrak{g}$. Consider arbitrary $x,y,z,t\in\mathfrak{g}$. Then
$\omega([\theta(x,y),z]+[\theta(y,z),x]+[\theta(z,x),y],t)=\omega(x,\theta(y,[z,t])+[t,\theta(y,z)]+\theta([y,t],z).$
Therefore, conditions (29) and (30) are equivalent. The proposition is proved.
Theorem 2. Let $(\mathfrak{g},\omega)$ be a quadratic finite-dimensional Lie algebra, where $\mathfrak{g}$ is a perfect Lie algebra with $Z(\mathfrak{g})=0$, $r\in\mathfrak{g}\wedge\mathfrak{g}$ be a skew-symmetric tensor, $R_{r}$ be the corresponding map, and $\theta:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$ be a bilinear map defined by (26) Then $(\mathfrak{g},\delta_{r})$ is a Lie bialgebra if and only if $\theta(x,y)=\nu([x,y])$ for some $\nu\in\mathrm{Cent}(\mathfrak{g})$.
Proof. If $\theta(x,y)=\nu([x,y])$ for some $\nu\in\mathrm{Cent}(\mathfrak{g})$, then $(\mathfrak{g},\delta_{r})$ is a Lie bialgebra by Proposition 4.
Suppose that $(\mathfrak{g},\delta_{r})$ is a Lie bialgebra. By Proposition 4 and Proposition 5, the map $\theta$ satisfies equations (17)-(19). By Statement 5, there is a map $\nu\in\mathrm{Cent}(\mathfrak{g})$ such that $\theta(x,y)=\nu([x,y])$ for all $x,y\in\mathfrak{g}$. The theorem is proved.
Theorem 3. Let $(\mathfrak{g},\omega)$ be a quadratic finite-dimensional Lie algebra, where $\mathfrak{g}$ is a perfect Lie algebra with $Z(\mathfrak{g})=0$, $r\in\mathfrak{g}\wedge\mathfrak{g}$ be a skew-symmetric tensor, $R_{r}$ be the corresponding map, and $\theta:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$ be a bilinear map defined by (26) Then
(i) The pair $(\mathfrak{g},\delta_{r})$ is a triangular Lie bialgebra if and only if $\theta(x,y)=0$ for all $x,y\in\mathfrak{g}$.
(ii) The pair $(\mathfrak{g},\delta_{r})$ is a quasitriangular Lie bialgebra if and only if $\theta(x,y)=-\mu^{2}([x,y])$ for some $\mu\in\mathrm{Cent}(\mathfrak{g})$ and all $x,y\in\mathfrak{g}$.
(iii) The pair $(\mathfrak{g},\delta_{r})$ is a factorizable Lie bialgebra if and only if $\theta(x,y)=-\mu^{2}([x,y])$ for some non-degenerate $\mu\in\mathrm{Cent}(\mathfrak{g})$ and all $x,y\in\mathfrak{g}$.
Proof. (i) This is equivalent to the fact that skew-symmetric solutions of the classical Yang-Baxter equations are in one-to-one correspondence with skew-symmetric Rota-Baxter operators of weight 0.
(ii) Suppose that $\theta(x,y)=-\mu^{2}([x,y])$ for some $\mu\in\mathrm{Cent}(\mathfrak{g})$.
Consider a map $B_{r}$ defined as
$B_{r}(x)=\frac{1}{2}(R_{r}(x)-\mu(x))$
for all $x\in\mathfrak{g}$. By Statement 2, $B_{r}$ is a Rota-Baxter operator of weight $\mu$. Since $R_{r}+R_{r}^{*}=0$ and $\mu^{*}=\mu$ (since $\mathfrak{g}$ is perfect), $B_{r}$ satisfies $B_{r}+B_{r}^{*}+\mu=0$. Let $r_{1}\in\mathfrak{g}\otimes\mathfrak{g}$ be the corresponding to the map $2B_{r}$ tensor (12). By Theorem 1, $(\mathfrak{g},\delta_{r_{1}})$ is a quasitriangular Lie bialgebra.
By Proposition 3, the multiplication in the dual algebra of the bialgebra $(\mathfrak{g},\delta_{r})$ is given by
$[a^{*},b^{*}]=(-[R(a),b]-[a,R(b)])^{*}=(-[2B_{r}(a),b]-[a,2B_{r}(b)]-2\mu[a,b])^{*}$
for all $a,b\in\mathfrak{g}$. That is, the multiplication in the dual algebra of the coalgebra $(\mathfrak{g},\delta_{r})$ coincides with the product in the dual algebra of the coalgebra $(\mathfrak{g},\delta_{r_{1}})$. Therefore, $\delta_{r}=\delta_{r_{1}}$ and $(\mathfrak{g},\delta_{r})$ is a quasitriangular Lie bialgebra.
Suppose that $(\mathfrak{g},\delta_{r})$ is a quasitriangular Lie bialgebra. That is, $\delta_{r}=\delta_{t}$, where $t\in\mathfrak{g}\otimes\mathfrak{g}$ is a solution of the CYBE and $t+\tau(t)$ is non-degenerate. By Theorem 1, the corresponding to $t$ map $R_{t}$ is a Rota-Baxter operator of weight $\eta\in\mathrm{Cent}(\mathfrak{g})$ and for all $x\in\mathfrak{g}$
$R_{t}(x)+R^{*}_{t}(x)+\eta(x)=0.$
Let $q=t-r=\sum\limits_{i}x_{i}\otimes y_{i}\in\mathfrak{g}\otimes\mathfrak{g}$. Since for any $x\in\mathfrak{g}$ $\delta_{r}(x)=\delta_{t}(x)$, we have that $[q,x]=0$ for any $x\in\mathfrak{g}$. Let $\mu:\mathfrak{g}\rightarrow\mathfrak{g}$ be the corresponding to $q$ map defined by (12). We have for any $y\in\mathfrak{g}$
$0=(y^{*}\otimes id)([q,x])=\sum\limits\omega(y,[x_{i},x])y_{i}+\omega(y,x_{i})[y_{i},x]=\mu([x,y])+[\mu(y),x].$
That is, $\mu\in\mathrm{Cent}(\mathfrak{g})$. Since $\mathfrak{g}$ is perfect, $\mu^{*}=\mu$. Therefore,
$R_{t}+R_{t}^{*}=R_{r}+R_{r}^{*}+2\mu=2\mu,$
that is, $\eta=-2\mu$. Let $x,y\in\mathfrak{g}$. Since $R_{t}=R_{r}+\mu$ is a Rota-Baxter operator of weight $\eta$, for all $x,y\in\mathfrak{g}$ we have
$0=[R_{t}(x),R_{t}(y)]-R_{t}([R_{t}(x),y]+[x,R_{t}(y)]+\eta([x,y]))\\ =[(R_{r}(x)+\mu(x),R_{r}(y)+\mu(y)]-R_{r}([R_{r}(x)+\mu(x),y]+[x,R_{r}(y)+\mu(y)]+\eta([x,y]))\\ -\mu([R(x)+\mu(x),y]+[x,R_{r}(y)+\mu(y)]+\eta([xy])\\ =[R_{r}(x),R_{r}(y)]-R_{r}([R_{r}(x),y]+[x,R_{r}(y)])+\mu^{2}([x,y]).$
Therefore, $\theta(x,y)=-\mu^{2}([x,y])$. The theorem is proved.
5. Simple finite-dimensional Lie bialgebras.
All the results in this section are not new. Nevertheless, it would be convenient to have a unique approach to the topic.
Definition. An (arbitrary) algebra $A$ over a field $F$ is called absolutely simple (normally simple in [23]) if any field extension $F\subset\overline{F}$, the algebra $A\otimes_{F}\overline{F}$ is simple.
We will need the following classical characterization of absolutely simple algebras.
Theorem 4[23]. An arbitrary finite-dimensional algebra $A$ is absolutely simple if and only if it is simple and $\mathrm{Cent}(\mathfrak{g})=\{\lambda{\rm id}|\ \lambda\in F\}$.
It is known that the complexification of $sl_{n}(\mathbb{R})$ is equal to $sl_{n}(\mathbb{C})$, that is, $sl_{2}(\mathbb{R})$ is absolutely-simple. As an example of a not absolutely simple Lie algebra, one can consider the special complex Lie algebra $\mathfrak{g}=sl_{2}(\mathbb{C})$ and consider $\mathfrak{g}$ as an algebra over $\mathbb{R}$ (there is a non-scalar centrum map in $\mathrm{Cent}(\mathfrak{g})$, see Example 2). In this case, the complexification $\overline{\mathfrak{g}}=\mathfrak{g}_{\mathbb{R}}\otimes\mathbb{C}$ is a semisimple Lie algebra ([19]).
The following theorem is a corollary of Theorem 3 (wich was originally proved in [2] for $F=\mathbb{R}$).
Theorem 5. Let $\mathfrak{g}$ be a simple finite-dimensional Lie algebra over a field $F$ of characteristic 0. Then any Lie bialgebra structure on $\mathfrak{g}$ is either triangular, factorizable, or almost-factorizable.
Proof. Let $\delta:\mathfrak{g}\rightarrow\mathfrak{g}\otimes\mathfrak{g}$ be a Lie bialgebra structure on $\mathfrak{g}$. By the first Whitehead’s Lemma, any Lie bialgebra structure $\delta$ on $\mathfrak{g}$ is cobaundary, that is, there exists $r\in\mathfrak{g}\wedge\mathfrak{g}$ such that $\delta=\delta_{r}$.
Let $R_{r}$ be the corresponding to $r$ map defined by (12) and $\theta$ be the corresponding bilinear map defined by (26). By Corollary 4, there is a map $\nu\in\mathrm{Cent}(\mathfrak{g})$ such that
$\theta(x,y)=\nu([x,y])$
for all $x,y\in\mathfrak{g}$. Since the kernel of $\nu$ is an ideal of $\mathfrak{g}$, either $\nu=0$ or $\nu$ is non-degenerate. If $\nu=0$, then by Theorem 3, $\delta_{r}$ is a triangular Lie bialgebra.
[…]
Since $\mathfrak{g}$ is a simple finite-dimensional Lie algebra, $\mathfrak{g}_{\bar{F}}$ is a finite-dimensional semisimple Lie algebra over an algebraically closed field $\overline{F}$. Let $\mathfrak{g}_{\overline{F}}=\bigoplus_{k=1}^{s}I_{k}$, where $I_{j}$ are simple ideals of $\mathfrak{g}_{\overline{F}}$. Obviously, for any $j\in\{1,\ldots,s\}$, $\overline{\nu}(I_{j})=I_{j}$ and the restriction $\overline{\nu}_{j}=\overline{\nu}|_{I_{j}}$ is a centrum map defined on $I_{j}$. By Schur’s lemma, $\overline{\nu}_{j}=\lambda_{j}{\rm id}$ for some $\lambda_{j}\in\overline{F}$, $\lambda_{j}\neq 0$ Since $\overline{F}$ is algebraically closed, there are $\alpha_{i}\in\overline{F}$ such that $\alpha_{i}^{2}=-\lambda_{i}$. Then the map $\mu$ acting as
$\mu(x_{1}+x_{2}+\ldots+x_{s})=\alpha_{1}x_{1}+\alpha_{2}x_{2}+\ldots+\alpha_{s}x_{s},$
$x_{j}\in I_{j}$ is a centrum map satisfying $\mu^{2}=-\overline{\nu}$. Thus, $(\mathfrak{g}_{\overline{F}},\overline{\delta}_{r})$ is a factorizable Lie bialgebra by Theorem 3, and $(\mathfrak{g},\delta)$ is an almost-factorizable Lie bialgebra.
Proposition 6. Let $\mathfrak{g}$ be a finite-dimensional simple, not absolutely simple, real Lie algebra. Then any Lie bialgebra on $\mathfrak{g}$ is either triangular or factorizable.
Proof. Consider $\mathrm{Cent}(\mathfrak{g})$. Since the kernel of any centrum map is an ideal in $\mathfrak{g}$, any nonzero element from $\mathrm{Cent}(\mathfrak{g})$ is invertible. Therefore, $\mathrm{Cent}(\mathfrak{g})$ is a field containing $\mathbb{R}$ as a proper subfield. Since $\mathrm{Cent}(\mathfrak{g})\subset\mathrm{End}(\mathfrak{g})$ and $\mathfrak{g}$ is a finite-dimensional algebra, $\mathrm{Cent}(\mathfrak{g})$ is a finite extension of $\mathbb{R}$. Thus, $\mathrm{Cent}(\mathfrak{g})$ is isomorphic to $\mathbb{C}$, and for any $\nu\in\mathrm{Cent}(\mathfrak{g})$, there exists $\mu\in\mathrm{Cent}(\mathfrak{g})$ such that $\mu^{2}=\nu$. By Theorem 3, any Lie bialgebra on $\mathfrak{g}$ is either triangular, or factorizable. The proposition is proved.
Example 6. Let $\mathfrak{g}=sl_{2}(\mathrm{R})$ with the standard basis $x,h,y$ (see Example 2). Consider $r=x\otimes y-y\otimes x$. From [18], we know that up to the action of $\mathrm{Aut}(\mathfrak{g})$, there are three non-isomorphic Lie bialgebra structures given by $r_{1}=h\wedge x$ (triangular), $r_{2}=x\wedge y$ (factorizable), and $r_{3}=h\wedge(x+y)$ (almost-factorizable).
Let $\omega$ be the bilinear form on $\mathfrak{g}$ defined by $\omega(x,y)=1$, $\omega(h,h)=2$. For the illustration of the presented theory, consider corresponding to $r_{i}$ linear maps $R_{i}$ (12) and bilinear maps $\theta_{i}$ (26).
We have that $R_{1}(x)=0$, $R_{1}(y)=-h$ and $R_{1}(h)=2x$. Then
$\theta_{1}(h,y)=[R_{1}(h),R_{1}(y)]-R_{1}([R_{1}(h),y]+[h,R_{1}(y)])=4x-R(2h)=0.$
Similarly, $\theta_{1}(x,h)=\theta_{1}(x,y)=0$. Thus, $(\mathfrak{g},\delta_{r_{1}})$ is a triangular Lie bialgebra by Theorem 3.
Consider $r_{2}$. We have that $R_{2}(x)=-x$, $R_{2}(y)=y$, $R_{2}(h)=0$. Then
$\theta_{2}(x,y)=[R_{2}(x),R_{2}(y)]-R_{2}([R_{2}(x),y]+[x,R_{2}(y)])=-h-R_{2}(h-h)=-h=-[x,y],$
$\theta_{2}(x,h)=[R_{2}(x),R_{2}(h)]-R_{2}([R_{2}(x),h]+[x,R_{2}(h)])=-R_{2}(2x)=2x=-[x,h].$
Similarly, $\theta_{2}(y,h)=-[y,h]$. Thus, $\theta_{2}(a,b)=-[a,b]$ for all $a,b\in\mathfrak{g}$ and $(\mathfrak{g},\delta_{2})$ is a factorizable Lie bialgebra by Theorem 3.
Consider $r_{3}$. In this case, $R_{3}(x)=R_{3}(y)=-h$, $R_{3}(h)=2(x+y)$. We have
$\theta_{3}(x,y)=[R_{3}(x),R_{3}(y)]-R_{3}([R_{3}(x),y]+[x,R_{3}(y)])=-R_{3}(2y+2x)=4h=4[x,y],$
$\theta_{3}(x,h)=-4x+4y-R_{3}(2h)=-8x=4[x,h],$
$\theta_{3}(y,h)=-4x+4y-R_{3}(-2h)=8y=4[y,h].$
Thus, $\theta_{3}(a,b)=4[a,b]$ for all $a,b\in\mathfrak{g}$. Thus, $(\mathfrak{g},\delta_{r_{3}})$ is a Lie bialgebra. Since $\mathfrak{g}$ is absolutely simple, $\mathrm{Cent}(\mathfrak{g})$ is isomorphic to $\mathbb{R}$. By Theorem 3, $(\mathfrak{g},\delta_{r_{3}})$ can’t be a triangular or a factorizable Lie bialgebra. By Proposition 5, $(\mathfrak{g},\delta_{r_{3}})$ is an almost-factorizable Lie bialgebra.
Example 7 Let $\mathfrak{g}=sl_{2}(\mathbb{R})\oplus sl_{2}(\mathbb{R})$ be the simple Lie algebra from Example 2. Then $\mathfrak{g}$ is a 6-dimensional simple, not absolutely simple, real Lie algebra with the basis $\{x,h,y,ix,ih,iy\}$, where $\{x,h,y\}$ is the standard basis of $sl_{2}(\mathbb{R})$ ([19]). Let $\omega$ be a bilinear map defined as
$1=\omega(x,y)=\frac{1}{2}\omega(h,h)=-\omega(ix,iy)=-\frac{1}{2}\omega(ih,ih).$
Simple computations show that $(\mathfrak{g},\omega)$ is a quadratic Lie algebra. Consider an element $r=h\wedge(x+y)-(i\otimes i)(h\wedge(x+y))$, the corresponding map $R$ defined as
$R(x)=R(y)=-h,\ R(h)=2(x+y),\ R(ix)=R(iy)=-ih,\ R(ih)=2i(x+y).$
Note that we have the following equation: $R(ia)=iR(a)$ for all $a\in\mathfrak{g}$. From the previous example, we obtain that the corresponding bilinear map $\theta$ (defined by (26)) satisfies
$\theta(a,b)=4[a,b]=-(2\varphi_{i})^{2}([a,b])$
for all $a,b\in\mathfrak{g}$, where $\varphi_{i}$ is the centrum map from Example 2 defined as $\varphi_{i}(u+iv)=iu-v$ for all $u,v\in sl_{2}(\mathbb{R})$. Thus, $(\mathfrak{g},\delta_{r})$ is a factorizable Lie bialgebra. In order to find the corresponding solution of CYBE, consider the map $B=\frac{1}{2}(R-2\varphi_{i})$. We have that
$B(x)=-\frac{1}{2}h-ix=-iB(ix),\ B(y)=-\frac{1}{2}h-iy=-iB(iy),$
$B(h)=x+y-ih=-iB(ih).$
By Statement 2, $B$ is a Rota-Baxter operator of weight $2\varphi_{i}$. Since $R_{r}$ is skew-symmetric, $B$ satisfies $B+B^{*}+2\varphi_{i}=0$. By Theorem 1, the corresponding to $B$ tensor
$r_{B}=-(\mathrm{id}\otimes\varphi_{i}+\varphi_{i}\otimes\mathrm{id})(\frac{1}{2}h\otimes h+x\otimes y+y\otimes x)+\frac{1}{2}(\mathrm{id}\otimes\mathrm{id}-\varphi_{i}\otimes\varphi_{i})(h\wedge(x+y)).$
is a solution of CYBE on $\mathfrak{g}$ with $\mathfrak{g}$-invariant symmetric part.
Proposition 7. Let $\mathfrak{g}$ be a simple finite-dimensional Lie algebra over a field of characteristic 0, $r\in\mathfrak{g}\wedge\mathfrak{g}$, and $R_{r}\in\mathrm{End}(\mathfrak{g})$ be the corresponding to $r$ map defined by (12). Then
(1) $(\mathfrak{g},\delta_{r})$ is a triangular Lie bialgebra if and only if the classical double $D(\mathfrak{g})$ is not a semisimple Lie algebra. In this case, $D(\mathfrak{g})=A\oplus D$ (a direct sum of vector spaces), where $D$ is the radical of $D(\mathfrak{g})$ such that $D^{2}=0$.
(2) $(\mathfrak{g},\delta_{r})$ is a factorizable Lie bialgebra if and only if $D(A)$ is a semisimple not simple Lie bialgebra isomorphic to $\mathfrak{g}\oplus\mathfrak{g}$.
(3) $(\mathfrak{g},\delta_{r})$ is an almost-factorizable Lie bialgebra if and only if $D(\mathfrak{g})$ is a simple, not absolutely simple, Lie algebra.
Proof. Suppose that $(\mathfrak{g},\delta_{r})$ is a Lie bialgebra. Let $D(\mathfrak{g})$ be the classical double of $(\mathfrak{g},\delta_{r})$. By Proposition 3, $D(\mathfrak{g})=D_{R_{r},0}(\mathfrak{g})$. Let $I$ be a proper ideal in $D(\mathfrak{g})$. By Proposition 1, $\mathrm{dim}(I)=\mathrm{dim}(\mathfrak{g})$. Recall that $D(\mathfrak{g})$ is a quadratic Lie algebra with the nondegenerate bilinear form $Q$ defined as $Q(x+f,y+g)=f(y)+g(x)$ for all $x,y\in\mathfrak{g}$, $f,g\in\mathfrak{g}^{*}$. Let $I^{\perp}$ be the orthogonal complement of $I$ in $D(\mathfrak{g})$ with respect to the form $Q$. Since $(D(\mathfrak{g}),Q)$ is a quadratic Lie algebra and $I$ is a proper ideal, $I^{\perp}$ is a proper ideal in $D(\mathfrak{g})$. Therefore, $\mathrm{dim}(I^{\perp})=\mathrm{dim}(\mathfrak{g})$. The intersection $I_{1}=I\cap I^{\perp}$ is also a proper ideal in $D(\mathfrak{g})$. Therefore, we have two options: $I=I^{\perp}$ or $I\cap I^{\perp}=0$.
If $I=I^{\perp}$, then $Q(I^{2},D(\mathfrak{g}))\subset Q(I,[I,D(\mathfrak{g})])=0$. Thus, $I^{2}=0$ and $(\mathfrak{g},\delta_{r})$ is a triangular Lie bialgebra (see, for example, [17]).
If $I\cap I^{\perp}=0$, then $I^{2}\neq 0$, $(I^{\perp})^{2}\neq 0$, and $D(\mathfrak{g})=I\oplus I^{\perp}$. Since any proper ideal of $I$ or $I^{\perp}$ is an ideal in $D(\mathfrak{g})$, $I$ and $I^{\perp}$ are simple Lie algebras. Therefore, $D(\mathfrak{g})$ is a sum of two simple ideals, and $(\mathfrak{g},\delta_{r})$ is a factorizable Lie bialgebra (see, for example, [15]). Now we can use Corollary 5 to complete the proof.
We want to finish with the following result
Theorem 6. Let $\mathfrak{g}$ be a finite-dimensional semisimple Lie algebra over an algebraically closed field of characteristic 0, $\omega:\mathfrak{g}\times\mathfrak{g}\rightarrow F$ is a nondegenerate invariant biliner form on $\mathfrak{g}$, and $\delta:\mathfrak{g}\rightarrow\mathfrak{g}\wedge\mathfrak{g}$ is a linear map. Then $(\mathfrak{g},\delta)$ is a Lie bialgebra if and only if $\delta=\delta_{r}$ for some $r\in\mathfrak{g}\otimes\mathfrak{g}$ such that the corresponding linear map $R_{r}$ defined by (12) is a Rota-Baxter operator of a weight $\mu\in\mathrm{Cent}(\mathfrak{g})$ satisfying
$R_{r}+R_{r}^{*}+\mu=0.$
Proof. Suppose that $(\mathfrak{g},\delta)$ is a Lie bialgebra. By the first Whitehead’s Lemma, any Lie bialgebra structure $\delta$ on $\mathfrak{g}$ is cobaundary, that is, there exists $q\in\mathfrak{g}\wedge\mathfrak{g}$ such that $\delta=\delta_{q}$. Let $R_{q}$ and $\theta:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$ be the maps defined by (12) and (26). By Theorem 2, $\theta(x,y)=\mu([x,y])$ for all $x,y\in\mathfrak{g}$, where $\mu\in\mathrm{Cent}(\mathfrak{g})$. Using a similar technique as in Theorem 5, we can find a centrum map $\nu\in\mathrm{Cent}(\mathfrak{g})$ such that $\nu^{2}=-\mu$. By Theorem 3, $(\mathfrak{g},\delta_{r})$ is either triangular or quasitriangular Lie bialgebra. By Theorem 1, $\delta_{q}=\delta_{r}$, where $r\in\mathfrak{g}\otimes\mathfrak{g}$ and the corresponding map $R_{r}$ is a Rota-Baxter operator of a weight $\mu\in\mathrm{Cent}(\mathfrak{g})$ satisfying (15).
The inverse statement follows from Theorem 1. The theorem is proved.